Article #630: Lustre D filesystem unavailable
Update - 2:25pm, 12/16/2013 The LustreD scratch filesystem has been returned to service and both the filesystem and scheduler appear to be working pro...
Update - 2:25pm, 12/16/2013 The LustreD scratch filesystem has been returned to service and both the filesystem and scheduler appear to be working pro...
All ITaP Research Computing systems are currently experiencing an issue with accessing network filesystems. A case has been opened with our vendor as...
The Fortress HPSS Archive is offline due to issues with their storage systems relating to the power loss on the West Lafayette campus in the wake of t...
Nearly all major clusters operated by ITaP Research Computing are stopped due to issues with their storage systems relating to the power loss on the W...
Update: 11:00pm, Nov. 12, 2013 ITaP storage engineers have returned the offline hardware to production and LustreC is back in production. Queues on Ha...
In the evening of 10/10/2013, the fileserver providing the "scratch96" filesystem serving some users of the Steele and Radon clusters suffer...
Update - 10:15 am Fortress is back in full production. Original Message: As of 8:00am, Thursday, September 19, the Fortress HPSS is temporarily unavai...
Update: May 13, 2013 11:00pm: LustreC has been returned to service. Carter, Hansen, and Peregrine1 are back in production with queues enabled. Update...
Resolved: As of about 4:45pm ET, the connectivity issue affecting the Fortress archive has been resolved. The HPSS archive is back in full production....
On April 24, 2013, network engineers will be relocating fiber optics that connect the Peregrine1 cluster to infrastructure in West Lafayette. This out...
Update: 8:12pm Scheduling on Carter has been resumed, and Carter is back in full production. Original Message: Beginning the morning of April 16, a nu...
Update: ITaP engineers have corrected the issue affecting the LustreC filesystem. The system is back in production. Job scheduling on Carter, Hansen a...
As of 9:00am, are seeing a problem with the LustreC scratch filesystem that serves Carter, Hansen, and Peregrine1. To prevent any more jobs from runn...
Update: As of about 11:00 am, the problem with the chilled water has been corrected, and scheduling has resumed on all RCAC clusters. Thank you for yo...
Campus chilled water serving the MATH data center is experiencing above-normal temperatures, and as a precaution, scheduling on the Coates, Rossmann,...
Update: Noon, 1/8/13 The power issue in MATH has been resolved. Power has been restored to the nodes in the Coates-A subcluster affected by the outage...
During scheduled network maintenance on network equipment connecting storage to ITaP clusters, all scheduling will be paused from 4-6pm. Running jobs...
Update: 10:00pm Tuesday As of 8:30pm Tuesday 21 August 2012, the LustreB filesystem has been returned to full service. Our storage engineers with assi...
Update - April 11, 2012 240pm At around 240pm, ITaP engineers have restored communications between the HPSS system and the tape library. Access to For...
Update : 1:45pm As of As of 1:45pm this afternoon, systems staff have completed patching the samba servers used to access storage systems. You should...